# Emergent Response Planning in LLM

### v1

@@ 6.3 Future Research Directions
Several key research directions emerge from our findings. Understanding the causal relationship between planning representations and generation outcomes stands as a primary challenge - while our probing results show these representations exist, determining whether and how they influence the generation process requires further investigation. Also, exploring whether similar planning phenomena emerge in multilingual or multimodal contexts could provide insights into how these capabilities develop across different domains and training objectives. Addressing these questions could require developing robust evaluation frameworks and more sophisticated probing techniques.
@@ Acknowledgements
We would like to thank Yuyu Fan, Jiachen Ma and anonymous reviewers for their valuable feedback and helpful discussions.
@@
Zhanhui Zhou proposed the emergent response planning in discussion with Zhichen Dong; proposed experimental tasks; made substantial writing contributions to abstract, introduction, and Section 3.

### v2

@@ 6.4 Future Research Directions
Several key research directions emerge from our findings: (1) Causal mechanisms of planning: Research could investigate whether and how planning representations directly influence token generation (e.g., via causal intervention experiments). Establishing causality is crucial for reliably leveraging these representations and understanding LLM decision-making. (2) Leverage planning for generation control: Future work might explore methods to detect and utilize pre-generation attribute predictions (e.g., key content points) for real-time steering. This could enable more efficient and precise control than post-hoc correction, potentially reducing computational waste and errors by allowing early adjustments based on predicted response properties. (3) Planning in multimodal contexts: Exploration of whether similar planning phenomena emerge in multimodal (e.g., image+text) LLMs could be valuable. Such research may provide insights into the universality of emergent planning and how modality impacts the development of these cognitive capabilities. (4) Planning-aware training: Developing objectives that explicitly reward alignment between early-plan encodings and final outputs (e.g., via consistency losses) represents another avenue. This may enhance coherence in complex tasks by grounding generation in initial intent.
@@ Acknowledgements
We would like to thank Yuyu Fan, Jiachen Ma and anonymous reviewers for their valuable feedback and discussions.
@@
Zhanhui Zhou first pitched the research idea of emergent response planning to Zhichen Dong; proposed experimental tasks; made substantial writing contributions to abstract, introduction, and Section 3.